OpenStudy (anonymous):

find the partial differential equation of z=1/x *(f(x-ay)+f(x+ay)) where a is an arbitrary constant..

OpenStudy (anonymous):

here z=f(x,y)....i.e; z is a function in x and y..

OpenStudy (anonymous):

you mean\[z=\frac{1}{x}(f(x-ay)+f(x+ay))\] ?

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

answer......plz..:(

OpenStudy (anonymous):

\[z_{x}=−\frac{1}{x^2}(f(x−ay)+f(x+ay))+\frac{1}{x}(f'(x−ay)+f'(x+ay))\]

OpenStudy (anonymous):

\[z_{y}=\frac{1}{x}(-af'(x−ay)+af'(x+ay))\]

OpenStudy (anonymous):

k...next?

OpenStudy (anonymous):

\[z_{xx}=\frac{2}{x^3}(f'(x−ay)+f'(x+ay))-\frac{1}{x^2}(f''(x−ay)+f''(x+ay))+z_{x}\]

OpenStudy (anonymous):

\[z_{yy}=\frac{a^2}{x}(f''(x−ay)+f''(x+ay))\]

OpenStudy (anonymous):

k..

OpenStudy (anonymous):

\[f''(x−ay)+f''(x+ay)=\frac{1}{a^2} x z_{yy}\]

OpenStudy (anonymous):

hmn...i got it....thanq very much...@mukushla ...i can do it now...

OpenStudy (anonymous):

\[f'(x−ay)+f'(x+ay)=xz_{x}+z\]

OpenStudy (anonymous):

ok